# Class-Independent Increment: An Efficient Approach for Multi-label Class-Incremental Learning

### v1

@@ IV-C Ablation Study
TABLE VI: Sensitive study of hyper-parameter $\lambda$.
$\lambda$ 0 0.0001 0.001 0.01 0.1 1 10
Last Acc (%) 71.49 71.55 71.62 72.05 72.18 72.10 71.86
We finally conducted a sensitivity study on $\lambda$ using the same benchmark, where we changed the value of $\lambda$ in the reasonable range from $0.0001$ to $10$. From Table VI, it can be observed that as $\lambda$ increases, the final mAP(%) gradually improves and reaches the highest accuracy 72.18% at 0.1. However, further increases in $\lambda$ indicate that excessively strong regularization constraints can impact other loss functions.
TABLE VII: Compare different methods on MS-COCO with different numbers of exemplars $M_{per}$ per class.
@@
Ours Last Acc 71.4 72.6 73.4 74.5 75.1
Avg Acc 78.1 78.6 79.1 79.4 79.8
The Influence of Buffer Size: Incremental Learning methods typically utilize additional memory space to store a fixed number of exemplars per old class, denoted as ($M_{per}$), to mitigate catastrophic forgetting. Intuitively, retaining more exemplars may enhance recognition performance but also come with a greater memory burden. Therefore, it is valuable to compare our approach against the rehearsal-based methods PodNet [23] and the SOTA method KRT [11] with different numbers of exemplars in the MS-COCO with B0-C10 setting (refer to Table VII).
We observe that our method consistently outperforms PodNet* and KRT, achieving a final accuracy of 74.5%. As the number of exemplars ($M_{per}$) decreases, the advantage of our method continues to grow. Specifically, with an ample buffer of $M_{per}=30$, our method outperforms KRT by 3.6% in final accuracy. Moreover, in the scenario with no old examples ($M_{per}=0$), our method exhibits a remarkable improvement of 5.5%. This implies that CLIN does not rely on the storage of old examples. It is noteworthy that when the $M_{per}=0$, the accuracy of the rehearsal-based method is seriously attenuated and our method significantly outperforms PodNet* by $\xrightarrow{}$71.4%) on last accuracy.
@@ V Conclusion
Limitation and Feature work. This paper explores and addresses the task of MLCIL from the perspective of incremental learning and has achieved promising results. In terms of label correlation, it attempts to use a Transformer decoder to extract local discriminative features for different labels and establish co-occurrence relationships between categories. However, it does not delve into the contribution of label dependencies in preventing forgetting and learning new classes. Moreover, due to the insensitivity to algorithm time complexity in incremental learning tasks, our paper does not discuss it.
In the future, we will further (1) investigate the co-occurrence relationship between labels in the query ((token))-to-label structure, using label dependencies as knowledge retention to prevent forgetting and to enhance multi-label classification performance. (2) Investigate the impact of unlearned potential categories in images on the task. (3) Extend the approach to more network frameworks, such as large-scale natural language models like CLIP and other multi-label application scenarios.

### v2

@@ 4.3 Ablation Study
Sensitive Study of Hyper-parameter: The hyper-parameters $\alpha$ and $\beta$ control the strength of $\mathcal{L}_{mc}$ and $\mathcal{L}_{kd}$ to together control the total optimization objective.
@@
  Ours   Last Acc   71.7   72.9   73.8   75.0   75.3
  Avg Acc   78.4   78.9   79.3   79.8   80.0
Table 8: Comparison of different methods on MS-COCO with various numbers of exemplars $M_{per}$ per class.
@@
The Influence of Buffer Size: Incremental Learning methods typically utilize additional memory space to store a fixed number of exemplars per old class, denoted as ($M_{per}$), to mitigate catastrophic forgetting. Intuitively, retaining more exemplars may enhance recognition performance but also come with a greater memory burden. Therefore, it is valuable to compare our approach against the rehearsal-based methods PodNet [6] and the SOTA method KRT [5] with different numbers of exemplars in the MS-COCO with B0-C10 setting (refer to Table 8).
[…]
@@ 5 Conclusion
Limitation and Feature work. This paper explores and addresses the task of MLCIL from the perspective of incremental learning and has achieved promising results. In terms of label correlation, it attempts to use a Transformer decoder to extract local discriminative features for different labels and establish co-occurrence relationships between categories.
However, it does not delve into the contribution of label dependencies in preventing forgetting and learning new classes. Moreover, due to the insensitivity to algorithm time complexity in incremental learning tasks, our paper does not discuss it.
In the future, we will further (1) investigate the co-occurrence relationship between labels in the query ((token))-to-label structure, using label dependencies as knowledge retention to prevent forgetting and to enhance multi-label classification performance. (2) Investigate the impact of unlearned potential categories in images on the task. (3) Explore the generalizability of our feature decoupling strategy across diverse datasets, particularly analyzing the performance nuances on fine-grained and small-scale benchmarks like PASCAL VOC, and subsequently extend the approach to more network frameworks, such as large-scale natural language models like CLIP, and other multi-label application scenarios.